A statistics professor gave an exam on performing the T-Test with σ unknown. The exam had the bonus question "Guess the average score on this final exam. If a T-Test is performed and the mean guess for the class is not significantly different from the true average exam score, everyone in the class will then receive two bonus points." There were 10 students that took the final exam. The true average exam score was 89.3. The professor conducts a one-mean hypothesis at the 5% significance level, to test if the mean guess for the class is not significantly different from the true average exam score of 89.3. For this test: H0:μ=89.3; Ha:μ≠89.3, which is a two-tailed test. Student guesses are given below. Use a TI-83, TI-83 Plus, or TI-84 calculator to test whether the students should receive the two bonus points. Identify the test statistic, t, and p-value from the calculator output, rounding to two decimal places.

92 85 84 70 80 85 99 72 91 82
Answer #1

#R version 3.3.2 x=c(92,85,84,70,80,85,99,72,91,82) length (x) mean(x) sd(x) Run it (Shift+Enter) Save it Put on a wall [+] Sn:10, S 8.82 To +est Ho: 89.3 Ha: Test statistic, 84-89.3 8.82/VT0 - 5. 3 At df = p(Itl>-1-90): Þ - value = 0. 0899 = 0. 0899